Analytical Geometry

find the value of k so that straight line 2x+3y+4+k(6x-y+12)= 0 is perpendicular to the line 7x+5y-4 =0

krishnan s
15 Years agoGrade

1 Answer

xyz xz
14 Years ago

   2x+3y+4+k(6x-y+12) = 0

=x(2-6k) + y(3-k) + (4+12k)=0

 slope is  (6k-2)/(3-k)

but this line is perpendicular to 7x+5y-4=0 and its slope is -7/5

imples

    (-7/5)(6k-2)/(3-k) = -1

implies k = 29/47